OpenStudy (anonymous):

Can someone help me simplify the expression (6+6i)-(2+i)

OpenStudy (anonymous):

The choices are a. 4i b. 4i+4 c. 4+5i d. 9i

OpenStudy (mathstudent55):

First, drop the first set of parentheses since they are not needed.

OpenStudy (anonymous):

6+6i-(2+i)

OpenStudy (anonymous):

it would be 4+5i right?

OpenStudy (mathstudent55):

(6 + 6i) - (2 + i) = = 6 + 6i - (2 + i) Good, you got it. Now you have a negative sign outside the remaining set of parentheses. That is the same as having a -1 there. See below: = 6 + 6i - 1(2 + i) You need to distribute the -1 by the parentheses. You multiply each term inside the parentheses by -1. What that ends up doing is changing each sign inside the parentheses. = 6 + 6i - 2 - i Ok so far?

OpenStudy (anonymous):

Yeah I got it! I understand this one. @mathstudent55

OpenStudy (mathstudent55):

Now just collect like terms. Add 6 and -2 Add 6i and -1

OpenStudy (mathstudent55):

= 6 + 6i - 2 - i = 4 + 5i